Operations Report – April 21th

Crew 280 Operations Report 21-04-2023

SOL: 5
Name of person filing report: Neus Sabaté (Mission Specialist) and Cesca Cufí (Crew Engineer))
Non-nominal systems:
Robotic observatory.
Rover Spirit.
Notes on non-nominal systems sections:
Replacement parts of Robotic observatory are scheduled to be installed during crew 261 rotation. Corroded battery wire in Rover Spirit. Replacement pending (Sergii). Rover can only be used for short distance EVA.
Spirit rover used: yes
Hours: 217.0
Beginning charge: 100%
Ending charge: 65%
Currently charging: yes
Opportunity rover used: no
Hours: 121.1
Beginning charge: –
Ending charge: –
Currently charging: no
Curiosity rover used: no
Hours: 230.1
Beginning charge: 100%
Ending charge: 65%
Currently charging: yes
Perseverance rover used: yes
Hours: 261.5
Beginning charge: –
Ending charge: –
Currently charging: no
General notes on rovers: none
Summary of Hab operations:
Water use: 22 gal (from 6PM Sol4 to 6PM Sol5)
Water (static tank): 398 gal (at 6:00PM)
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: no
Summary of internet: used from 18:00 to 21:00 for personal communication and reports.
Summary of suits and radios:
Suits: 3, 6, 7 and 8.
Radios: EVA radios work nominally. Channel 1 com with the station was lost after passing by the road passage at Kissing Camel ridge, driving South. The EVA-Station com was maintained thanks to the crew’s own radio equipment (Midland HAM radios).
All radio and spacesuit batteries checked before EVA: ok.
Summary of GreenHab operations:
Water use: 14 gal for watering and 1 gal for experiments in Science Dome
Heater: off
Supplemental light: off
Harvest: 289 g cucumber (1 piece), 36 g micro greens, 163 g tomatoes, 1 g chives, 27g jazz mix, 1g salvia and 1g thyme
Neus Sabaté has taken 200gr of soil from Green Hab wheelbarrow to grow seeds for her Martian Batteries experiment at the Science Dome.
Summary of Science Dome operations:
Laia Ribas (Green Hab Officer): experiment.
Neus Sabaté (Mission Specialist): experiment.
Cesca Cufí (Crew Engineer): experiment
Dual split: off
Summary of RAM operations: CO monitor has been replaced.
Summary of any observatory issues:
Robotic observatory: Not working. Replacement parts are scheduled to be installed during crew 261 rotation.
Summary of health and safety issues:
Summary of other system’s issues: non
Questions, concerns and requests to Mission Support:
Today’s EVA has taken the team to Robert’s Garden Rock, which is relatively close to the base (approx. 1 mile) and the rover’s batteries discharged to 65%. We have observed that the batteries show a fast discharging rate at the beginning of the EVA and then keep discharging at a lower pace. Considering the Mission Support advice of not letting the rover’s batteries go down below 60% we would like to know if it is possible to plan a longer EVA that allows us to reach longer distances. Advice about this matter will be highly appreciated.
Spirit rover: replacement of the corroded wire
Kitchen: knobs replacement
Retrieval of trash in the rear airlock

Copyright © The Mars Society. All rights reserved. | Main Site